Consider the energy associated with the “Big Mac” meal:   
Big Mac 540 Calories
Medium fries: 340 Calories
Medium Coke: 220 Calories

If you have the Big Mac meal for lunch, how many (i) Calories (ii) calories (iii) Joules have you consumed?

Suppose you try to “lose” the calories gained when eating the “Big Mac meal” by exercising – say by doing an equivalent amount of work in the gymnasium by lifting a 50 kg “weight” a vertical distance of 5 feet (convert to metre).

- To compensate for the calories gained by eating the Big Mac meal, how much total work in Joules must be expended by exercising ?
- How much work is expended in lifting the 50 kg weight 5 feet once?
- How many times must you raise the “weight” to expend the total energy gained from food?
Answer #1

Total calories in big mac meal = 540+340+220 = 1100 calories (a answer)

b

1 cal = 4.184J

Thus 1100 cal = 4.184*1100 =4602.4J

Total work need to expended by exercise is 4602.4J

Energy used in once=mgh = 50*9.8*(5*0.305m) = 747.25J

Total times need to raise =4602.4/747.25 = 6.16

That's after 7 times all energy will be expend of it